Find a counterexample to the following statement:
If p is a prime number, then 2p-1 is prime.

Answer by rapaljer(4671)

How about 5, 13, or 23? There are a TON of examples of primes p such 2p-1 is NOT a prime.

5*2-1 = 9 which is NOT a prime.
2*13 - 1 = 25 which is NOT a prime.
2*23 - 1 = 45 which is NOT a prime.

R^2 at SCC